Chemist has 20% and 50% solution of acid available. how many liters of each solution should be mixed to obtain 10 liters of 40% acid solution?
Answer by jorel1380(3719)

x+y=10
.2x+.5y=4
2x+5y=40
2x+2y=20
3y=20
y=20/3
x=10/3
You need 10/3 ltrs. of 20% solution, and 20/3 ltrs. of 50% solution..